if ax²+bx+c is a quadratic equation. IF two persons A and B are there.if A person wrote the constant term c, wrong, he got roots as 8,2. IF B person wrote the coefficient of x² wrong he got roots as -9,-1.then find the actual quadratic equation and their roots?

Grade:12

1 Answers

Nishant Vora IIT Patna
7 years ago
Hey

The quadratic eqn for person A,
(x-2)(x-8)=0
(x^2 - 10x + 16)=0------------------(1)
Comparing this with the standard form
x^2 + \frac{b}{a}x + \frac{c}{a} = 0
we get,
b/a = -10

Now the quadratic Eqn for person B is :-
(x+9)(x+1)=0
(x^2 + 10x + 9)=0
x^2 + \frac{b}{a}x + \frac{c}{a} = 0
In this the leadingcoefficient is wrong
Comparing this with the standard form
we get, c/a = 9


so b = -10a and c= 9a

so QE will be
ax^2 -10 ax + 9a = 0
x^2 -10 x + 9 = 0